# Provably Efficient Online RLHF with One-Pass Reward Modeling

### v2

@@ Abstract
Reinforcement Learning from Human Feedback (RLHF) has shown remarkable success in aligning Large Language Models (LLMs) with human preferences. Traditional RLHF approaches rely on a fixed dataset, which often suffers from limited coverage. To this end, online RLHF has emerged as a promising direction, enabling iterative data collection and model improvement. Despite its potential, this paradigm faces a key bottleneck: the requirement to continuously integrate new data into the historical dataset and re-optimize the model from scratch at each iteration, resulting in computational and storage costs that grow linearly with the number of iterations. In this work, we address this challenge by proposing a one-pass reward modeling method that does not require storing the historical data and can be computed in constant time. Specifically, we first formalize RLHF as a contextual preference bandit problem and design an online mirror descent algorithm with a tailored local norm to replace the standard maximum likelihood estimation for reward modeling. We then apply our method to various online RLHF settings, including passive data collection, active data collection, and deployment-time adaptation. We provide theoretical guarantees showing that our method improves both statistical and computational efficiency. Finally, we provide practical algorithms and conduct experiments using Llama-3-8B-Instruct and Qwen2.5-7B-Instruct models on the Ultrafeedback-binarized and Mixture2 datasets, validating the effectiveness of our proposed method.
@@ 1 Introduction
Traditional RLHF methods primarily rely on fixed preference datasets, which typically suffer from limited coverage. As a result, the learned reward models struggle to generalize to out-of-distribution (OOD) samples, constraining the effectiveness of the aligned models. To address this, online RLHF has emerged as a promising paradigm, enabling iterative data collection and model improvement. The general process can be described as (i) collect the preference data; (ii) update the model using the collected preference data. The above two steps are repeated for several iterations to boost model performance. In practice, the Claude (Bai et al., 2022) and LLaMA-2 (Touvron et al., 2023) projects have demonstrated that online RLHF can significantly enhance model performance.
Despite its empirical success, online RLHF introduces significant computational challenges. Specifically, the typical process of online RLHF involves continuously integrating newly collected data into the dataset and re-optimizing the model from scratch over the expanded dataset. While this strategy is statistically efficient, its computational and storage costs scale linearly with the number of iterations, which becomes prohibitive in long-term iterations, especially on edge devices where computation and memory resources are inherently limited. This raises a pressing question:
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
In this work, we provide an affirmative answer to this question in the setting of contextual preference bandits with linearly parameterized reward functions. Specifically, building on recent theoretical advancements in RLHF (Zhu et al., 2023; Das et al., 2024; Ji et al., 2024), we formulate the RLHF problem as a contextual dueling bandit problem (Yue et al., 2012; Saha, 2021). While prior work has explored this formulation, most existing methods focus on statistical efficiency and overlook the growing computational burden. To bridge this gap, we introduce a novel one-pass reward modeling algorithm which does not require storing the historical data and can be computed in constant time. Our method is based on the online mirror descent framework with a tailored local norm that captures second-order information. We then apply our method to several online RLHF settings, including passive data collection, active data collection, and deployment-time adaptation. We establish theoretical guarantees showing that our method improves both statistical and computational efficiency. Table 1 summarizes the comparison of our method with the existing methods.
To enable usage in LLMs, we develop practical variants of our method. Direct computation and storage of the Hessian matrix is prohibitively expensive; thus, we propose an efficient approximation using Hessian-Vector Products (HVP) combined with conjugate gradient descent, avoiding explicit second-order information and relying only on first-order computation. Additionally, we employ rejection sampling to approximate model uncertainty in a computationally efficient manner. With the above techniques, we conduct experiments using the LLaMA-3-8B-Instruct (Llama Team, 2023) and Qwen2.5-7B-Instruct (Qwen Team, 2024) models on the Ultrafeedback-binarized (Cui et al., 2023) and Mixture2 (Dong et al., 2024) datasets, validating the effectiveness of our method.
@@
  • We apply our method to a broad range of online RLHF settings, including passive data collection, active data collection, and deployment-time adaptation. For each setting, we design tailored algorithms and establish corresponding theoretical guarantees, demonstrating that our approach achieves improved statistical and computational efficiency than existing methods.
  • We develop practical algorithms by approximating the update using Hessian-Vector Products combined with conjugate gradient descent, and estimating uncertainty via rejection sampling. Based on the above techniques, we conduct empirical evaluations using LLaMA-3-8B-Instruct and Qwen2.5-7B-Instruct models on Ultrafeedback-binarized and Mixture2 datasets, showing our method achieves better performance in both statistical and computational efficiency.
Organization. The rest of the paper is organized as follows: Section 3 introduces the problem setup. Section 4 presents our proposed one-pass reward modeling method. Section 5 applies our method to different online RLHF settings. Section 6 provides the practical implementations. Section 7 presents experimental results. Section 8 concludes the paper. We defer the proofs, experiment details, and more experimental results to the appendix.
@@ 2 Related Work
In this section, we review the works most closely related to ours, including offline RLHF, online RLHF, contextual dueling bandits, and active learning.
Offline RLHF.
Learning from human preferences with deep learning models dates back to Christiano et al. (2017) and has been recently popularized by the success of large language models (Bai et al., 2022; Touvron et al., 2023; Achiam et al., 2023). Existing results generally fall into two categories: reward-based and reward-free. Reward-based methods typically consist of two steps: reward modeling and reward maximization (Ouyang et al., 2022). Given preference data, the reward model is trained to predict preference labels, and the LLM is fine-tuned based on the reward model using RL algorithms such as PPO (Schulman et al., 2017). In contrast, reward-free methods directly optimize the LLM using human preference feedback, without relying on explicit reward modeling (Rafailov et al., 2023; Azar et al., 2024). Among these works, Direct Preference Optimization (DPO) (Rafailov et al., 2023) is one of the most popular reward-free methods, which treats generative models directly as reward models and optimizes them using human preference feedback. Given the empirical success of RLHF, recent efforts have been devoted to developing a deeper theoretical understanding of this approach. Zhu et al. (2023) proposed to formulate RLHF as a contextual preference bandit problem and proved the maximum likelihood estimator converges under both the Bradley-Terry (BT) model and Plackett-Luce (PL) model with linear reward functions.
@@
Offline RLHF methods rely on fixed preference datasets, which often suffer from limited data coverage. Consequently, the resulting reward models struggle to generalize to out-of-distribution samples, thereby limiting the effectiveness of the aligned models. To overcome this limitation,online RLHF has emerged as a promising alternative, enabling iterative data collection and continuous model refinement. The works (Dong et al., 2023; Guo et al., 2024; Yuan et al., 2024; Wu et al., 2025) have demonstrated that online iterative variants of direct preference learning algorithms significantly outperform their offline counterparts. Xiong et al. (2024) identified key challenges in offline RLHF and theoretically demonstrated the potential benefits of online exploration. Recent work has incorporated optimism-driven bonus terms into the DPO loss to encourage exploration in online RLHF (Xie et al., 2025; Cen et al., 2025; Zhang et al., 2025). These studies primarily focus on the sample efficiency, but do not consider the accompanying increase in computational complexity. In this paper, we propose algorithms that achieve both statistical and computational efficiency.
@@
Dueling bandits are a variant of the multi-armed bandit problem in which the learner sequentially selects a pair of arms and receives noisy binary feedback (Yue et al., 2012). The contextual dueling bandit framework extends this setting by incorporating contextual information (Dudík et al., 2015; Saha, 2021; Bengs et al., 2022). Within this framework, Saha (2021) studied the $K$-armed contextual dueling bandit problem, and Saha et al. (2023) further extended it to the reinforcement learning setting. Additionally, Sekhari et al. (2024) investigated the contextual dueling bandit problem under an active learning paradigm, where the learner adaptively queries for feedback, aiming to minimize both regret and the number of queries. To move beyond linear reward functions, Verma et al. (2025) introduced the neural dueling bandit problem, modeling the reward function using neural networks. These prior works commonly rely on maximum likelihood estimation to learn the reward function, leading to computational complexity that grows linearly with the number of iterations. In contrast, we propose algorithms that maintain constant per-iteration computational complexity, while preserving statistical efficiency.
@@
Active learning is a paradigm that aims to reduce the labeling cost by selecting the most informative samples for annotation (Settles, 2009). In general, existing work can be categorized into two settings: pool-based and stream-based. The pool-based setting (Seung et al., 1992; Freund et al., 1997; Huang et al., 2010) involves the learner iteratively selecting a batch of informative samples from a large unlabeled pool, querying their labels, updating the model, and repeating this process. In contrast, the stream-based setting (Cesa-Bianchi et al., 2004, 2006; Cacciarelli and Kulahci, 2024) requires the learner to sequentially observe data points and decide in real time whether to query their labels. Within the context of RLHF, Das et al. (2024) studied pool-based active learning, while Ji et al. (2024) focused on the stream-based setting. In this work, we focus on the pool-based strategy, which can be naturally extended to the stream-based scenario.
@@ 3 Problem Setup
Following recent advancements in RLHF (Zhu et al., 2023; Das et al., 2024; Xiong et al., 2024), we formulate RLHF as a contextual bandit problem. Specifically, we have a set of contexts $\mathcal{X}$ and a set of possible actions $\mathcal{A}$ per context. To learn with human preference feedback, the learner selects a tuple $(x,a,a^{\prime})$ to present to the human, where $x\in\mathcal{X}$ is the context, $a,a^{\prime}\in\mathcal{A}$ are the actions. The human then provides a binary preference feedback $y\in\{0,1\}$, where $y=1$ indicates that the human prefers action $a$ over action $a^{\prime}$, and $y=0$ otherwise. We study the commonly used Bradley-Terry (BT) model in preference learning (Bradley and Terry, 1952), which assumes that the human’s preference is generated by a logistic function of the difference in the rewards of the two actions.
@@
Given a context $x\in\mathcal{X}$ and two actions $a,a^{\prime}\in\mathcal{A}$, the probability of the human preferring action $a$ over action $a^{\prime}$ is given by
$\mathbb{P}\left[y=1\mid x,a,a^{\prime}\right]=\frac{\exp\left(r(x,a)\right)}{\exp\left(r(x,a)\right)+\exp\left(r\left(x,a^{\prime}\right)\right)},$ (1)
where $r:\mathcal{X}\times\mathcal{A}\to\mathbb{R}$ is a latent reward function.
We consider the widely used linear reward in the literature (Zhu et al., 2023; Cen et al., 2025).
@@
The feature mapping $\phi$ can be constructed by removing the last layer of a pre-trained large language model, and $\theta^{*}$ corresponds to the weights of the last layer.
@@
The non-linearity coefficient $\kappa$ is defined as
$\kappa=\max_{x\in\mathcal{X},a,a^{\prime}\in\mathcal{A}}\max_{\theta\in\Theta}\frac{1}{\dot{\sigma}\left(\phi(x,a)^{\top}\theta-\phi(x,a^{\prime})^{\top}\theta\right)},$ (2)
where $\dot{\sigma}(w)=\sigma(w)(1-\sigma(w))$ is the derivative function.
The quantity $\kappa$ captures the learning complexity that satisfies $\kappa\leq 2+\exp(2BL)+\exp(-2BL)$. It may be exceedingly large with exponential dependence on the scale of features and parameters.
@@ 4.2 One-pass Reward Modeling
where ¯$\bar{\mathcal{H}}_{t}=\sum_{i=1}^{t-1}H_{i}\left(\bar{\theta}_{i+1}\right)+\lambda I$ is the local norm, and $\eta$ is the step size. The optimization problem can be decomposed into two terms. The first term is the instantaneous log-loss $\ell_{t}(\theta)$, which accounts for the information of the current sample. The second consists of a quadratic proxy for the past losses constructed through the sequence $\{\bar{\theta}_{i}\}_{i\leq t}$. A key component is the design of the local norm ¯$\bar{\mathcal{H}}_{t}$, which approximates the Hessian matrix by $H_{i}\left(\bar{\theta}_{i+1}\right)$ at a lookahead point $\bar{\theta}_{i+1}$. Such a Hessian matrix effectively captures local information and is crucial for ensuring statistical efficiency.
@@
Standard OMD.  To enhance computational efficiency, a natural alternative is to replace this formulation with the standard OMD framework, which permits a closed-form solution and thus eliminates the need for iterative optimization. However, the standard OMD minimizes a first-order approximation of the loss function, which sacrifices several key properties compared to its implicit counterpart, as demonstrated by Campolongo and Orabona (2020). Specifically, the standard OMD formulation updates using $g_{t}(\theta_{t})$, whereas the implicit OMD updates the algorithm approximately with the subsequent sub-gradient, $g_{t}(\theta_{t+1})$. This distinction results in a notable gap in the convergence rates of the two methods. To this end, we propose to approximate the current loss $\ell_{t}(\theta)$ using a second-order Taylor expansion, drawing inspiration from Zhang and Sugiyama (2023). Define the second-order approximation of $\ell_{t}(\theta)$ as ~$\widetilde{\ell}_{t}(\theta)=\ell_{t}(\widetilde{\theta}_{t})+g_{t}(\widetilde{\theta}_{t})^{\top}(\theta-\widetilde{\theta}_{t})+\frac{1}{2}\lVert\theta-\widetilde{\theta}_{t}\rVert_{H_{t}(\widetilde{\theta}_{t})}^{2}$. Then, we replace the loss $\ell_{t}(\theta)$ in Eq. (4) with the approximation ~$\widetilde{\ell}_{t}(\theta)$, leading to the update rule:
@@
where $\eta$ is the step size and ~$\widetilde{\mathcal{H}}_{t}=\mathcal{H}_{t}+\eta H_{t}(\widetilde{\theta}_{t})$ is the local norm with $\mathcal{H}_{t}\triangleq\sum_{i=1}^{t-1}H_{i}(\widetilde{\theta}_{i+1})+\lambda I$. Eq. (5) can be solved with a projected gradient step with the following equivalent form:
@@
Thus, the estimator $\widetilde{\theta}_{t+1}$ provides a closed-form solution, leading to a $\mathcal{O}(1)$ computational complexity per iteration. Since the estimator processes the samples in a one-pass manner, it mitigates the memory burden associated with computing the gradient of the full dataset. These properties make the method particularly suitable for edge devices, where both memory and computational resources are severely constrained. The detailed process of our proposed method is presented in Algorithm 1.
@@
2:  Update ~$\widetilde{\mathcal{H}}_{t}=\mathcal{H}_{t}+\eta H_{t}(\widetilde{\theta}_{t})$
@@
0:  Regularization $\lambda$, step size $\eta$
1:  Initialize $\widetilde{\theta}_{1}=\bm{0}$ and ~$\widetilde{\mathcal{H}}_{1}=\lambda I$

### v3

@@ Abstract
Reinforcement Learning from Human Feedback (RLHF) has shown remarkable success in aligning Large Language Models (LLMs) with human preferences. Traditional RLHF methods rely on a fixed dataset, which often suffers from limited coverage. To this end, online RLHF has emerged as a promising direction, enabling iterative data collection and refinement. Despite its potential, this paradigm faces a key bottleneck: the requirement to continuously integrate new data into the dataset and re-optimize the model from scratch at each iteration, resulting in computational and storage costs that grow linearly with the number of iterations. In this work, we address this challenge by proposing a one-pass reward modeling method that eliminates the need to store historical data and achieves constant-time updates per iteration. Specifically, we first formalize RLHF as a contextual preference bandit and develop a new algorithm based on online mirror descent with a tailored local norm, replacing the standard maximum likelihood estimation for reward modeling. We then apply it to various online RLHF settings, including passive data collection, active data collection, and deployment-time adaptation. We provide theoretical guarantees showing that our method enhances both statistical and computational efficiency. Finally, we design practical algorithms for LLMs and conduct experiments with the Llama-3-8B-Instruct and Qwen2.5-7B-Instruct models on Ultrafeedback and Mixture2 datasets, validating the effectiveness of our approach.
@@ 1 Introduction
Traditional RLHF methods primarily rely on fixed preference datasets, which typically suffer from limited coverage. As a result, the learned reward models struggle to generalize to out-of-distribution samples, constraining the effectiveness of the aligned models. To address this, online RLHF has emerged as a promising paradigm, enabling iterative data collection and model improvement. The general process can be described as (i) collect the preference data; (ii) update the model using the collected data. The above two steps are repeated for several iterations to boost model performance. In practice, Claude (Bai et al., 2022) and LLaMA-2 (Touvron et al., 2023) have demonstrated that online RLHF can significantly enhance model performance (Dong et al., 2024). Theoretically, recent works (Xie et al., 2025; Cen et al., 2025) indicate that online exploration can improve the statistical efficiency of RLHF. Beyond performance gains, online RLHF serves as a crucial step toward agentic applications, where models can continuously integrate environmental feedback to enable real-time interaction, adaptive reasoning, and autonomous decision-making (Silver and Sutton, 2025).
Despite its empirical success, online RLHF may introduces significant computational challenges. Specifically, the typical process of online RLHF involves continuously integrating newly collected data into the historical dataset and re-optimizing the model from scratch over the expanded dataset. While this strategy is statistically efficient, its computational and storage costs scale linearly with the number of iterations, which becomes prohibitive in long-term iterations, especially on edge devices where computation and memory resources are inherently limited. This raises a pressing question:
@@
Table 1: Comparison between previous works and our work in terms of the statistical and computational efficiency across different online RLHF settings. The column “Context” and “Action” represent the context and action are determined by the environment () or the algorithm (). For the computational efficiency (time and storage), we highlight the dependence on the $t$ at iteration $t$. Here, $d$ is the feature dimension, $T$ is the total number of iterations, $\kappa$ is the non-linearity coefficient, $\Phi=\mathbb{E}_{x\sim\rho}\left[\phi(x,\pi^{*}(x))\right]$ is the concentrability vector, $V_{T}$ and $\mathcal{H}_{T}$ are two local norms satisfying $\lVert\Phi\rVert_{\mathcal{H}_{T}^{-1}}\leq\sqrt{\kappa}\lVert\Phi\rVert_{V_{T}^{-1}}$ (*: amortized complexity over $T$).
@@
Passive $\widetilde{\mathcal{O}}\big(\sqrt{d}\cdot\kappa\lVert\Phi\rVert_{V_{T}^{-1}}\big)$ $\mathcal{O}(\log T)^{*}$ $\mathcal{O}(t)$ Zhu et al. (2023)
@@
Active $\widetilde{\mathcal{O}}\big(d\sqrt{\kappa/T}\big)$ $\mathcal{O}(t\log t)$ $\mathcal{O}(t)$ Das et al. (2025)
@@
Deploy $\widetilde{\mathcal{O}}\big(d\kappa\sqrt{T}\big)$ $\mathcal{O}(t\log t)$ $\mathcal{O}(t)$ Saha et al. (2023)
@@
In this work, we provide an affirmative answer to this question in the setting of contextual preference bandits with linearly parameterized reward functions. Specifically, building on recent theoretical advancements in RLHF (Zhu et al., 2023; Das et al., 2025; Ji et al., 2025), we formulate the RLHF problem as a contextual dueling bandit problem (Yue et al., 2012; Saha, 2021). While prior work has explored this formulation, most existing methods focus on statistical efficiency and overlook the growing computational burden. To bridge this gap, inspired by recent progress in bandit learning (Zhang and Sugiyama, 2023; Li et al., 2024), we introduce a novel one-pass reward modeling method based on the online mirror descent framework with a tailored local norm that captures second-order information. Unlike traditional approaches, our method eliminates the need to store historical data and achieves constant-time updates per iteration, i.e., the computational cost remains invariant with respect to the cumulative number of iterations. We then apply our method to several online RLHF settings, including passive data collection, active data collection, and deployment-time adaptation. We establish theoretical guarantees showing that our method improves both statistical and computational efficiency. Table 1 summarizes the comparison of our method with the existing works.
To enable usage in LLMs, we develop practical variants of our method. Direct computation and storage of the Hessian matrix is prohibitively expensive; thus, we propose an efficient approximation using Hessian-Vector Products (HVP) combined with conjugate gradient descent, avoiding explicit second-order information and relying only on first-order computation. Additionally, we employ rejection sampling to approximate model uncertainty in a computationally efficient manner. With the above techniques, we conduct experiments using the LLaMA-3-8B-Instruct (Llama Team, 2023) and Qwen2.5-7B-Instruct (Qwen Team, 2024) models on the Ultrafeedback (Cui et al., 2024) and Mixture2 (Dong et al., 2024) datasets. Experimental results validate the effectiveness of our method.
@@
  • We apply our method to a broad range of online RLHF settings, including passive data collection, active data collection, and deployment-time adaptation. For each setting, we design tailored algorithms and establish corresponding theoretical guarantees, demonstrating that our approach achieves improved statistical and computational efficiency compared to existing methods.
  • We develop practical algorithms by approximating the update using Hessian-Vector Products combined with conjugate gradient descent, and estimating uncertainty via rejection sampling. Based on the above techniques, we conduct empirical evaluations using the LLaMA-3-8B-Instruct and Qwen2.5-7B-Instruct models on the Ultrafeedback and Mixture2 datasets, showing that our method improves both statistical and computational efficiency compared to existing methods.
Organization. Section 2 reviews the related work. Section 3 introduces the problem setup. Section 4 presents our proposed one-pass reward modeling method and section 5 applies it to various online RLHF settings. Section 6 provides practical versions of our method. Section 7 presents experimental results. Section 8 concludes the paper. The proofs and experiment details are deferred to the appendix.
@@ 2 Related Work
In this section, we review the works most closely related to ours, including online RLHF, contextual dueling bandits, and active learning.
@@
Traditional RLHF methods predominantly rely on fixed datasets, which often suffer from limited data coverage. Consequently, the resulting reward models struggle to generalize to out-of-distribution samples, thereby limiting the effectiveness of the aligned models. To overcome this limitation, online RLHF has emerged as a promising alternative, enabling iterative data collection and continuous model refinement. The works (Dong et al., 2023; Guo et al., 2024; Yuan et al., 2024; Wu et al., 2025) have demonstrated that online iterative variants of direct preference learning algorithms significantly outperform their offline counterparts. Xiong et al. (2024) identified key challenges in offline RLHF and theoretically demonstrated the potential benefits of online exploration. Recent work has incorporated optimism-driven bonus terms into the objective to encourage exploration in online RLHF (Xie et al., 2025; Cen et al., 2025; Zhang et al., 2025; Zhao et al., 2025). These approaches primarily focus on the sample efficiency, but do not consider the accompanying increase in computational complexity. To improve computational efficiency, Foster et al. (2025) tackled the challenge of enumerating an exponentially large response space. Differently, our work focuses on alleviating the computational burden that scales linearly with the number of iterations in online RLHF.
@@
Dueling bandits are a variant of the multi-armed bandit problem in which the learner sequentially selects a pair of arms and receives binary feedback (Yue et al., 2012). The contextual dueling bandit framework extends this setting by incorporating contextual information (Dudík et al., 2015; Saha, 2021; Bengs et al., 2022). Within this framework, Saha (2021) studied the $K$-armed contextual dueling bandit problem, and Saha et al. (2023) further extended it to the reinforcement learning setting. Additionally, Sekhari et al. (2023) investigated the contextual dueling bandit problem under an active learning paradigm, where the learner adaptively queries to minimize both regret and the number of queries. To move beyond linear reward functions, Verma et al. (2025a) introduced the neural dueling bandit problem, modeling the reward function using neural networks. These prior works commonly rely on maximum likelihood estimation to learn the reward function, leading to computational complexity that grows linearly with the number of iterations. In contrast, we propose algorithms that maintain constant per-iteration computational complexity.
@@
Active learning is a paradigm that aims to reduce the labeling cost by selecting the most informative samples for annotation (Settles, 2009). In general, existing work can be categorized into two settings: pool-based and stream-based. The pool-based setting (Seung et al., 1992; Freund et al., 1997; Huang et al., 2010) involves the learner iteratively selecting a batch of informative samples from a large unlabeled pool, querying their labels, updating the model, and repeating this process. In contrast, the stream-based setting (Cesa-Bianchi et al., 2004, 2006; Cacciarelli and Kulahci, 2024) requires the learner to sequentially observe data points and decide in real time whether to query their labels. Within the context of RLHF, Das et al. (2025) and Verma et al. (2025b) studied pool-based active learning, while Ji et al. (2025) focused on the stream-based setting. In this work, we focus on the pool-based strategy, which can be naturally extended to the stream-based scenario.
@@ 3 Problem Setup
Following recent advancements in RLHF (Zhu et al., 2023; Das et al., 2025; Xiong et al., 2024), we formulate RLHF as a contextual bandit problem. Specifically, we have a set of contexts $\mathcal{X}$ and a set of possible actions $\mathcal{A}$ per context. To learn with human preference feedback, the learner selects a tuple $(x,a,a^{\prime})$ to present to the human, where $x\in\mathcal{X}$ is the context, $a,a^{\prime}\in\mathcal{A}$ are the actions. The human then provides a binary preference feedback $y\in\{0,1\}$, where $y=1$ indicates that the human prefers action $a$ over action $a^{\prime}$, and $y=0$ otherwise. We study the commonly used Bradley-Terry (BT) model in preference learning (Bradley and Terry, 1952), which assumes that the human’s preference is generated by a logistic function of the difference in the rewards of the two actions.
@@
Given a context $x\in\mathcal{X}$ and two actions $a,a^{\prime}\in\mathcal{A}$, the probability of the human preferring action $a$ over action $a^{\prime}$ is given by $\mathbb{P}\left[y=1\mid x,a,a^{\prime}\right]=\frac{\exp\left(r(x,a)\right)}{\exp\left(r(x,a)\right)+\exp\left(r\left(x,a^{\prime}\right)\right)}$, where $r:\mathcal{X}\times\mathcal{A}\to\mathbb{R}$ is a latent reward function.
To facilitate theoretical analysis, following prior works (Zhu et al., 2023; Cen et al., 2025), we consider the linear realizable setting, where the reward function is parameterized by a linear model.
@@
While this setting is a simplification of the real-world problem, it serves as a useful and analytically tractable starting point. Specifically, the feature mapping $\phi$ can be obtained by removing the final layer of a pre-trained large language model, with $\theta^{*}$ corresponding to the weights of that layer. Moreover, this assumption can be further relaxed by allowing model misspecification (Jin et al., 2020) and neural function approximation (Du et al., 2024; Verma et al., 2025b).
@@
Let $\dot{\sigma}(w)=\sigma(w)(1-\sigma(w))$ be the derivative function of $\sigma$, the non-linearity coefficient $\kappa$ is defined as $\kappa=\max_{x\in\mathcal{X},a,a^{\prime}\in\mathcal{A},\theta\in\Theta}\frac{1}{\dot{\sigma}(\phi(x,a)^{\top}\theta-\phi(x,a^{\prime})^{\top}\theta)}$.
Intuitively, the quantity $\kappa$, defined as the inverse of the derivative, characterizes the learning difficulty of the reward function. In particular, a smaller derivative leads to a larger $\kappa$, implying that the model output changes less for the same input variation and thus the function is harder to learn. By direct calculation, we have $\kappa\leq 3+\exp(2BL)$. Therefore, $\kappa$ can be exceedingly large, exhibiting an exponential dependence on the magnitude of the features and the model parameters.
@@ 4.2 One-pass reward modeling
where $\bar{\mathcal{H}}_{t}=\sum_{i=1}^{t-1}H_{i}(\bar{\theta}_{i+1})+\lambda I$ is the local norm, and $\eta$ is the step size. The optimization problem can be decomposed into two terms. The first term is the instantaneous log-loss $\ell_{t}(\theta)$, which accounts for the information of the current sample. The second consists of a quadratic proxy for the past losses constructed through the sequence $\{\bar{\theta}_{i}\}_{i\leq t}$. A key component is the design of the local norm $\bar{\mathcal{H}}_{t}$, which approximates the Hessian matrix by $H_{i}(\bar{\theta}_{i+1})$ at a lookahead point $\bar{\theta}_{i+1}$. Such a Hessian matrix effectively captures local information and is crucial for ensuring statistical efficiency.
@@
Standard OMD.  To enhance computational efficiency, a natural alternative is to replace this formulation with the standard OMD framework, which permits a closed-form solution and thus eliminates the need for iterative optimization. However, the standard OMD minimizes a first-order approximation of the loss function, which sacrifices several key properties compared to its implicit counterpart, as demonstrated by Campolongo and Orabona (2020). Specifically, the standard OMD formulation updates using $g_{t}(\theta_{t})$, whereas the implicit OMD updates the algorithm approximately with the subsequent sub-gradient, $g_{t}(\theta_{t+1})$. This distinction results in a notable gap in the convergence rates of the two methods. To this end, we propose to approximate the current loss $\ell_{t}(\theta)$ using a second-order Taylor expansion, drawing inspiration from Zhang and Sugiyama (2023). Define the second-order approximation of $\ell_{t}(\theta)$ as $\widetilde{\ell}_{t}(\theta)=\ell_{t}(\widetilde{\theta}_{t})+g_{t}(\widetilde{\theta}_{t})^{\top}(\theta-\widetilde{\theta}_{t})+\frac{1}{2}\lVert\theta-\widetilde{\theta}_{t}\rVert_{H_{t}(\widetilde{\theta}_{t})}^{2}$. Then, we replace the loss $\ell_{t}(\theta)$ in Eq. (2) with the approximation $\widetilde{\ell}_{t}(\theta)$, leading to the update rule:
@@
where $\eta$ is the step size and $\widetilde{\mathcal{H}}_{t}=\mathcal{H}_{t}+\eta H_{t}(\widetilde{\theta}_{t})$ is the local norm with $\mathcal{H}_{t}\triangleq\sum_{i=1}^{t-1}H_{i}(\widetilde{\theta}_{i+1})+\lambda I$. Eq. (3) can be solved with a projected gradient step with the following equivalent form:
@@
Thus, the estimator $\widetilde{\theta}_{t+1}$ provides a closed-form solution, leading to a $\mathcal{O}(1)$ computational complexity per iteration. Furthermore, since the estimator processes the samples in a one-pass manner, it mitigates the memory burden associated with computing the gradient of the full dataset. These properties make the method particularly suitable for edge devices, where both memory and computational resources are severely constrained. The detailed process of our proposed method is presented in Algorithm 1.
@@
2: Update $\widetilde{\mathcal{H}}_{t}=\mathcal{H}_{t}+\eta H_{t}(\widetilde{\theta}_{t})$
@@
0: Regularization parameter $\lambda$, step size $\eta$
1: Initialize $\widetilde{\theta}_{1}=\bm{0}$ and $\widetilde{\mathcal{H}}_{1}=\lambda I$